# Discovering and exploiting active sensing motifs for estimation

### v1

@@ 7. Reveal active sensing motifs by correlating state dynamics with observability
$\mathbf{y}=\begin{bmatrix}r_{x}\\ \end{bmatrix}=\begin{bmatrix}v_{x}/z\\ \end{bmatrix}$, (12)
where the inputs $u_{z}$ and $u_{x}$ are measured accelerations. We chose this model for its simplicity and generality—it should be applicable to virtually any flying agent.
We designed a data-driven ANN estimator that can accurately estimate altitude given a two-second time window of optic flow and forward acceleration measurements (LABEL:fig:figure_supp_altitude_ANN_aLABEL:fig:figure_supp_altitude_ANN_b). We designed our ANN and training data to assume a constant altitude, thus we did not include the vertical acceleration as an input and left that for the KF to utilize. We augmented the measurement vector with the altitude estimate
@@
Guided by our observability analysis, which revealed that the variance of our altitude estimator should be strongly correlated with lateral acceleration (LABEL:fig:figure_supp_altitude_ANN_cLABEL:fig:figure_supp_altitude_ANN_d), we constructed a function for the altitude estimate’s noise variance $\check{\mathcal{R}}_{z,z}$ based on the empirical relationship between the mean lateral acceleration magnitude in the ANN time window (LABEL:fig:figure_supp_altitude_ANN_cLABEL:fig:figure_supp_altitude_ANN_d). To match the rapid decline of altitude estimation errors with increasing lateral accelerations we chose a function with a negative exponential form,
@@
The value of $\sigma$ can range between $[0,1]$, ensuring that for windows with lateral accelerations close to the maximum $\sigma=0$ and $\rho_{\min}$ will dominate, leading to a small value for $\check{\mathcal{R}}_{z,z}$, allowing the ANN altitude estimate to have a strong influence on the Kalman filter update equation. For our system, we chose $\rho_{\min}=10^{-3}$ to allow the accurate augmented altitude measurements have a large influence when acceleration was large, and we set $\rho_{\max}=10^{12}$ to ensure that periods of low acceleration did not affect the AI-UKF. For simplicity, we set the off-diagonal elements of $\check{\mathcal{R}}_{z}$ to 0. Although a correlation analysis of the ANN estimator may provide the most statistically consistent values for the off-diagonal elements, we go on to show that the AI-KF can be effective without this step.
To test the AI-UKF, we simulated a straight trajectory with a brief ($3\text{\,}\sec$) deceleration and then brief acceleration (LABEL:fig:figure_4_b). We added zero-mean Gaussian measurement noise for the entire trajectory length and a small amount of biased noise between the deceleration and acceleration (LABEL:fig:figure_4_c). The brief periods of high observability and unmodeled biased noise combined with nonlinear dynamics make this a prime example for demonstrating the utility of the AI-UKF.
Figure 5: Augmented Information Kalman Filter application for outdoor quadcopter flight. a. Quadcopter (DJI Matrice 300 RTK) flying in an outdoor environment with a representative ventral camera image and computed optic flow vectors. b. An experimental measured quadcopter trajectory with estimated altitude observability computed using the approach in Figure 1, illustrating how our BOUNDS framework can be used to determine observability for measured, rather than simulated, trajectories. c. Same as LABEL:fig:figure_4_c but for experimental data from the trajectory in b. As in LABEL:fig:figure_4_c, the clear benefit of the AI-UKF is its robustness to different initial conditions and faster convergence to correct estimates.
@@
We applied both a UKF and AI-UKF to the trajectory while varying the guess for the initial altitude $z_{0}$ (LABEL:fig:figure_4_d). We found that the performance of a regular UKF was overall poor (LABEL:fig:figure_4_d). For some initial conditions, the UKF altitude estimate moved in the correct direction during the deceleration (first period of high observability) (LABEL:fig:figure_4_d). However, the UKF had a hard time fully converging to the true altitude and correcting during the acceleration (second period of high observability) after being thrown off by the biased noise (LABEL:fig:figure_4_d). In contrast, the AI-UKF was able to converge quickly and smoothly to the true altitude during both periods of high observability (LABEL:fig:figure_4_d). Due to the coupling of altitude and forward velocity in the optic flow measurement ($r_{x}=v_{x}/z$), the forward velocity estimate also converged close to the true value for the AI-UKF, but again not for the UKF (LABEL:fig:figure_4_d). This illustrates how just a single data-driven estimator may be needed to estimate correlated state variables via the AI-KF.
@@
We also performed a sensitivity analysis to understand how the choice for the initial covariance $P_{k=0}$ and tuning of the process noise covariance matrix $Q$ affected the relative performance of the AI-UKF compared to the UKF. We specifically compared the median error after the first acceleration for the UKF and AI-UKF (LABEL:fig:figure_4_e). We found that the AI-UKF consistently outperformed the UKF (LABEL:fig:figure_4_e). While it may be possible to choose a specific combination of $x_{k=0}$, $P_{k=0}$, and $Q$ for a regular UKF that approximately matches the AI-UKF performance, there is not a universal choice that will work on all trajectories and initializations (LABEL:fig:figure_4_e). Therefore, the AI-UKF provides a degree of robustness to these initialization and hyperparameter choices.
@@
Another important consideration is that the amount of observability, or amount of acceleration/deceleration in this case, will influence the advantage of the AI-UKF. To test the generality of the AI-UKF, we applied it to a range of simulated trajectories with varying accelerations and found that the AI-UKF consistently matched or exceeded the performance of the UKF (LABEL:fig:figure_4_f). In general, for large enough accelerations/decelerations, both the UKF and AI-UKF performed equally well (LABEL:fig:figure_4_f). However, the UKF struggled with smaller accelerations/decelerations, while the AI-UKF was able to handle these instances without issue (LABEL:fig:figure_4_f). For very small or zero accelerations the UKF and AI-UKF were equivalent because there was not enough observability for the augmented data-driven state estimate to have an effect (LABEL:fig:figure_4_f). Overall, the AI-UKF appears to be most beneficial during periods of medium observability where a UKF might fail.
@@
Lastly, we confirmed that the AI-UKF performs well on real data. We collected optic flow and acceleration measurements from a quadcopter flying with approximately constant altitude and sporadic decelerations/accelerations, and then computed the observability of the trajectory via BOUNDS (LABEL:fig:figure_5_aLABEL:fig:figure_5_b). We found that we could use the same data-driven altitude estimator trained on simulated trajectories (Figure S6) for the real measurements, and it yielded impressive accuracy (LABEL:fig:figure_5_c). However, when applying a regular UKF we observed similar results to our simulations, wherein the filtered estimates would sometimes not converge for certain initializations (LABEL:fig:figure_5_c). In contrast, AI-UKF estimates converged quickly to the true altitude and forward velocity (LABEL:fig:figure_5_c). For some initializations the AI-UKF converged in seconds whereas the UKF did not converge within the $\sim$50-second trajectory (LABEL:fig:figure_5_c). Overall, we suggest that in situations with limited or sparse sensor data, like in GPS-denied environments, estimating mission-critical state variables can benefit greatly from the AI-KF framework.
@@ Discussion
We introduced a novel framework “BOUNDS: Bounding Observability for Uncertain Nonlinear Dynamic Systems” for empirically evaluating the observability of individual state variables. BOUNDS can be applied to discover active sensing motifs and design corresponding control strategies for autonomous systems or applied to evaluate data collected from moving organisms or machines. In most cases, active sensing motifs are likely to be turning or changes in ground speed, either of the body, or the sensors themselves. BOUNDS can help reveal exactly which combinations of these movements and measurements improve state estimation along observed trajectories of animals, or planned trajectories of machines. Implementing BOUNDS from scratch can be cumbersome, so we provide a Python package pybounds that automates these steps with numerous examples (see Code). We additionally introduced a harmonious framework for leveraging observability information about individual state variables—from BOUNDS—to intelligently filter state estimates. By integrating this idea with traditional Kalman filtering approaches, we showed how data-driven individual state estimates that are weighted by observability can be used to augment and improve state estimation with a framework we call the Augmented Information Kalman Filter. We employed a flying agent model to illustrate the applications of our novel methods, but our general framework is amenable to any system leveraging active sensing for estimation.
Tips and Tricks
BOUNDS has just two primary hyper-parameters, 1) the regularization parameter $\lambda$ (Eq. 3), and 2) the window size $\omega$ over which to analyze the observability. It is best practice to choose $\lambda$ to be large enough to numerically invert the Fisher information matrix, but small enough to not distort the total information. In general, a good guideline is to choose $\lambda$ to be less than the smallest nonzero eigenvalue of the Fisher information matrix Horn2012. Alternatively, one could symbolically compute the limit in Eq. 3, however, this can be computationally expensive and can yield infinite values corresponding to unobservable state variables. For $\omega$, we recommend choosing a window size roughly equal to the slowest motif of interest. For higher order systems it may be beneficial to take a more automated approach for discovering motifs using techniques either through methodical search cellini2024discovering, or genetic algorithms, particle swarm optimization, or reinforcement learning Kennedy. Other practical suggestions for applying our approach are provided in the Methods.

### v2

@@ Results
$\mathbf{y}=\begin{bmatrix}r_{x}\\ \end{bmatrix}=\begin{bmatrix}-v_{x}/z\\ \end{bmatrix}$, (7)
where the inputs $u_{z}$ and $u_{x}$ are measured accelerations. We chose this model for its simplicity and generality—it applies to virtually any flying agent.
We trained a data-driven ANN estimator that accurately estimates altitude from two-second windows of optic flow and forward acceleration (LABEL:fig:figure_supp_altitude_ANN_aLABEL:fig:figure_supp_altitude_ANN_b), and augmented the KF measurement vector with this altitude estimate,
@@
Guided by insight from our observability analysis (LABEL:fig:figure_supp_altitude_ANN_cLABEL:fig:figure_supp_altitude_ANN_d), we developed an empirical relationship between the altitude estimate’s noise variance $\check{\mathcal{R}}_{z,z}$ and the mean forward acceleration magnitude from the ANN time window (LABEL:fig:figure_supp_altitude_ANN_cLABEL:fig:figure_supp_altitude_ANN_d). To match the rapid decline of altitude estimation errors with increasing horizontal accelerations we chose a function with a negative exponential form,
@@
The value of $\sigma$ lies in the range $[0,1]$, ensuring that for windows with horizontal accelerations close to the maximum $\sigma=0$ and $\rho_{\min}$ will dominate, leading to a small value for $\check{\mathcal{R}}_{z,z}$, allowing the ANN altitude estimate to have a strong influence on the Kalman filter update equation. For our system, we chose $\rho_{\min}=10^{-3}$ to allow the accurate augmented altitude measurements have a large influence when acceleration was large, and we set $\rho_{\max}=10^{12}$ to ensure that periods of low acceleration did not affect the AI-UKF. For simplicity, we set the off-diagonal elements of $\check{\mathcal{R}}_{z}$ to 0.
To Test the AI-UKF, we simulated a straight trajectory with brief ($3\text{\,}\sec$) deceleration and acceleration (LABEL:fig:figure_4_b), zero-mean Gaussian noise throughout, and biased noise between maneuvers (LABEL:fig:figure_4_c). As expected, the ANN estimates accurately predicted the correct altitude only during periods of relatively high observability (LABEL:fig:figure_4_c).
Figure 5: Augmented Information Kalman Filter application for outdoor quadcopter flight. a. Quadcopter (DJI Matrice 300 RTK) flying in an outdoor environment with a representative ventral camera image and computed optic flow vectors. b. An experimental measured quadcopter trajectory with estimated altitude observability computed using the approach in Figure 1, illustrating how our BOUNDS framework can be used to determine observability for measured, rather than simulated, trajectories. c. Same as LABEL:fig:figure_4_c but for experimental data from the trajectory in b. As in LABEL:fig:figure_4_c, the clear benefit of the AI-UKF is its robustness to different initial conditions and faster convergence to correct estimates. Ground truth velocity and altitude was provided by GPS measurements.
@@
Applying both UKF and AI-UKF while varying initial altitude guesses (LABEL:fig:figure_4_d) showed poor UKF performance overall. For some initial conditions, the UKF altitude estimate moved in the correct direction during the deceleration (first period of high observability) (LABEL:fig:figure_4_d). However, the UKF struggled to converge to the true altitude after being thrown off by the biased noise (LABEL:fig:figure_4_d). In contrast, the AI-UKF was able to converge quickly and smoothly to the true altitude during both periods of high observability (LABEL:fig:figure_4_d). Due to the coupling of altitude and forward velocity in the optic flow measurement ($r_{x}=v_{x}/z$), the forward velocity estimate also converged close to the true value for the AI-UKF, but again not for the UKF (LABEL:fig:figure_4_d). This illustrates how just a single data-driven estimator may be needed to estimate correlated state variables via the AI-KF.
@@
In a sensitivity analysis, we examined how initial covariance $P_{k=0}$ and process noise $Q$ affected relative performance of the AI-UKF compared to the UKF. The AI-UKF consistently outperformed the UKF in terms of the median error after the first acceleration (LABEL:fig:figure_4_e). Although it may be possible to choose a specific combination of $\mathbf{x}_{k=0}$, $P_{k=0}$, and $Q$ for a regular UKF that approximately matches the AI-UKF performance, there is no universal choice that will work on all trajectories and initializations (LABEL:fig:figure_4_e). Therefore, the AI-UKF provides a degree of robustness to these initialization and hyperparameter choices.
@@
The observability level—here, acceleration magnitude—drives the AI-UKF advantage. For large enough accelerations/decelerations, both the UKF and AI-UKF performed equally well (LABEL:fig:figure_4_f). However, the UKF struggled with smaller accelerations/decelerations, while the AI-UKF was able to handle these instances without issue (LABEL:fig:figure_4_f). For very small or zero accelerations the UKF and AI-UKF were equivalent because there was not enough observability for the augmented data-driven state estimate to have an effect (LABEL:fig:figure_4_f). Overall, the AI-UKF is most beneficial during periods of medium observability where a UKF might fail.
@@
Fianlly, we validated the AI-UKF on real data. We collected optic flow and acceleration measurements from a quadcopter flying with approximately constant altitude and sporadic decelerations/accelerations, and then computed the observability of the trajectory via BOUNDS (LABEL:fig:figure_5_aLABEL:fig:figure_5_b). During these decelerations (LABEL:fig:figure_5_c) the quadcopter also changed course (LABEL:fig:figure_5_b). However, because our analysis is restricted to the velocity, optic flow, and acceleration along the direction of travel, these changes in course have no impact on the observability or estimates (LABEL:fig:figure_2_cv) . The same data-driven altitude estimator trained on simulated trajectories (Figure S6) yielded impressive accuracy on real measurements during periods of high observability (LABEL:fig:figure_5_d). However, when applying a regular UKF we observed similar results to our simulations, wherein the filtered estimates would sometimes not converge for certain initializations (LABEL:fig:figure_5_e). In contrast, AI-UKF estimates converged quickly to the true altitude and forward velocity (LABEL:fig:figure_5_e). For some initializations, the AI-UKF converged in seconds whereas the UKF did not converge within the $\sim$50-second trajectory (LABEL:fig:figure_5_c). Overall, we suggest that in situations with limited or sparse sensor data, like in GPS-denied environments, estimating mission-critical state variables can benefit greatly from the AI-KF framework.
@@ Discussion
Decoding and exploiting active sensing requires tools that bridge formal observability theory with real-world complexity. BOUNDS builds this bridge by empirically quantifying the observability of individual state variables along trajectories for nonlinear, partially observable systems. Drawing on these observability insights, the Augmented Information Kalman filter demonstrates how to statistically merge classic and data-driven estimates through an observability-weighted function. This provides stability during periods of poor observability while exploiting information from sporadic sensing maneuvers. By accessing a larger history of information, this approach enables recovery from poor initialization that would corrupt sequential nonlinear filters. Beyond methodology, our work also reveals a key active sensing principle: state variables often have distinct observability properties requiring tailored strategies for sensor selection and movement. For example, for flying agents—be they machines or organisms—wind direction estimation is best supported by changes in course, whereas estimating altitude can be achieved solely through straight line acceleration (Figure 2).
Our framework facilitates bidirectional insights for understanding natural navigation and advancing autonomous systems. In biology, a critical aspect of active sensing is knowing when to perform it. Many animals appear to balance dedicated sensing maneuvers with other navigation goals while considering energetic cost 16 through exploration-exploitation trade-offs triggered when estimation error reaches a threshold 4. BOUNDS provides a principled framework for generating related predictions: by computing observability along behavioral trajectories, we can quantify when specific state estimates will degrade and which movements would restore accuracy. For example, our analysis predicts that wind direction estimation degrades during straight flight but is restored by heading changes—testable via closed-loop experiments where heading turns are blocked or disrupted before flies turn upwind 53. Because neurons often multiplex sensory signals with complex filter dynamics 15, it can be difficult to understand which navigational variables can be extracted. BOUNDS can help by revealing which combinations of movement and neural activity patterns are sufficient for estimating specific state variables, generating testable hypotheses about computational roles without requiring explicit decoding 38.
How animals integrate sporadic information from active sensing bouts into continuous estimates remains an open question. Our observability filter and AI-KF provide a theoretical structure for exploring this. If neural circuits implement observability-weighted integration similar to the AI-KF, then state estimates should update rapidly during high-observability movements but remain stable during low-observability periods—a prediction testable through neural recordings during naturalistic behavior. Already some evidence supports observability-gated updates: Drosophila throttle learning plasticity in the central complex based on turning 22, a motif we find corresponds to high observability for many state variables (Fig. LABEL:fig:figure_2_c), suggesting that learning may be selectively enabled during observable epochs.
Model-based estimation remains standard in robotics 56, yet nonlinear tasks increasingly use windowed optimization 58, 48, 39 or data-driven methods 42—approaches that often sacrifice stability guarantees, simplicity, and interpretability 29. The AI-KF bridges these worlds through a statistically principled approach that is inherently modular: individual data-driven estimators (e.g., monocular SLAM, deep visual odometry) can be added or removed in the AI-KF structure without retraining. While not universally superior, the AI-KF excels for systems with sparse and dynamic observability that challenge traditional methods. Together, BOUNDS and the AI-KF enable engineers to develop machines that use active sensing to minimize required sensors, plan informative trajectories, and design robust estimators to approach the versatility observed in biological systems.
Methods
BOUNDS step by step detail
2. Define a model
$\mathbf{x}_{k+1}$ $=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k}),$ (11)
where $\mathbf{x}_{k}$ is the true state vector representing $n$ state variables, $\mathbf{u}_{k}$ is the input vector, and $\mathbf{\tilde{y}}_{k}$ is the measurement vector corresponding to $p$ measurements at time step $k$. The function $\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})$ describes the system dynamics. The function $\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$ describes the measurement dynamics, and $\mathbf{v}_{k}$ is zero-mean additive Gaussian noise with covariance matrix $R_{k}$. The effect of the noise on observability is taken care of in Step 5. Steps 3-4 only use the deterministic measurement equation $\mathbf{y}_{k}=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})$. A benefit of our empirical approach is that the model does not have to be closed-form, and can instead be a simulation, computational, or data-driven model 49, 31—for instance, a physics engine model of Drosophila 62 or finite element model of a flapping wing 7. The only requirement is the ability to simulate the system forward in time.
3. Reconstruct state trajectory through model predictive control
If control inputs are not known, such as for an observed trajectory, we apply model predictive control (MPC) 12 to find the inputs $\mathbf{U}$ required to reconstruct the state trajectory with the chosen model (LABEL:fig:figure_1_b). This step is critical for nonlinear or time-varying models where observability can change as a function of the state and input values. The MPC determined inputs are used to simulate the nominal trajectory $\mathbf{X}$ (which should be very similar to the observed trajectory $\mathbf{\tilde{X}}$) and associated measurements $\mathbf{Y}$ (LABEL:fig:figure_1_c).
4. Construct observability matrix
To quantify the observability of individual state variables, we next compute the observability matrix $\mathcal{O}$ empirically. $\mathcal{O}$ principally encodes information about the sensitivity of the sensor measurements $\mathbf{y}_{k}$ to small changes in an initial state $\mathbf{x}_{0}$. It can be compactly represented in discrete-time as the Jacobian of the measurements over time with respect to the initial state, $\Delta\mathbf{Y}_{\omega}/\Delta\mathbf{x}_{0}$ 24. Here we use $\mathbf{Y}_{\omega}$ to denote the collection of $p$ measurements for each time step from k=[0:w1]k=[0:w-1]. Following established methods49, 31, 14, we empirically approximate the Jacobian by applying small perturbations ($\epsilon=10^{-5}$) to each initial state variable $x_{i,0}$, one at a time, in the positive ${x}^{+}_{i,0}$ and negative ${x}^{-}_{i,0}$ directions. We simulate the system using the inputs $\mathbf{u}_{k}$ calculated in step 3 (they are not recalculated for the perturbed trajectory), and then compute the difference in the corresponding change in sensory measurements over time $\Delta\mathbf{Y}_{\omega}^{i}=\mathbf{Y}_{\omega}^{i+}-\mathbf{Y}_{\omega}^{i-}$ (LABEL:fig:figure_1_d). The end result for each state perturbation is a column vector of length $p\omega$, corresponding to $p$ measurements over $\omega$ time steps. We set a window size of $\omega=5$ steps for all simulations in this paper (see tips and tricks below for advice on selecting $\omega$). Repeating these steps for each state variable and stacking the resulting column vectors side-by-side yields the full empirically determined Jacobian, $\Delta\mathbf{Y}_{\omega}/\Delta\mathbf{x}_{0}=\Delta\mathbf{Y}_{\omega}/2\epsilon$, which is equivalent to the empirical observability matrix $\mathcal{O}$ (LABEL:fig:figure_1_e).
This empirical approach can lead to misleading values if the measurement function has any discontinuities, such as angular wrapping issues. For our simulations, all angular measurement variables were first unwrapped to ensure that $\Delta\mathbf{Y}$ did not have any discontinuities. An alternative for angular measurements is to replace them with the sine and cosine of each angular measurement.
5. Compute the Fisher information matrix and its inverse
To compute observability levels for individual state variables in physically interpretable units, we construct the Fisher information matrix (LABEL:fig:figure_1_f) as
$\mathcal{F}=\mathcal{O}^{\intercal}\mathcal{R}^{-1}\mathcal{O},$ (12)
where $\mathcal{R}$ is a block diagonal matrix of the measurement noise covariance matrices at each time step 8. In our simulations, we set $\mathcal{R}=0.1\cdot I_{(p\omega)\times(p\omega)}$, where $p$ is the number of measurements and $\omega$ is the sliding window size, assuming a noise variance of $0.1\text{\,}\mathrm{r}\mathrm{a}\mathrm{d}$ for angular measurements, $0.1\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ for airspeed and ground speed magnitude measurements, and $0.1\text{\,}{\mathrm{s}}^{-1}$ for optic flow measurements. The absolute noise properties of these measurements do not have a large impact on our results because larger or smaller values of $\mathcal{R}$ simply scale $\mathcal{F}$; however, relative changes in $\mathcal{R}$ for different measurements would have an effect. For the deterministic dynamics we consider, the process noise $Q$ is 0 (see the following section for incorporating nonzero $Q$). The Fisher information matrix captures the amount of information that a known random variable, in our case the measurements $\mathbf{Y}_{w}$, carries about unknown parameters, in our case the initial state $\mathbf{x}_{0}$. The Cramér–Rao bound states that the inverse of the Fisher information matrix $\mathcal{F}^{-1}$ sets the lower bound on the error covariance of the state estimate 19. Thus, the diagonal elements of $\mathcal{F}^{-1}$ tell us the minimum possible variance in the error of any unbiased estimate for each state variable, inversely correlating with observability (LABEL:fig:figure_1_g). In cases where there is at least one unobservable state variable, $\mathcal{F}$ will not be invertible. Therefore, to compute $\mathcal{F}^{-1}$ we add a small regularization term to ensure invertibility 17:
For unobservable state variables, this “Chernoff inverse” 17, 46 yields infinity in the corresponding diagonal elements, whereas the more traditional Moore-Penrose pseudoinverse 43 yields misleading values that are often small when they should be very large to provide a correct interpretation with the Cramér–Rao bound. Note that the regularization need not be limited to the form $\lambda I_{n\times n}$; a more general regularization of the form $\lambda J$, where $J$ is any $n\times n$ positive definite matrix, is also valid, though for simplicity we set $J=I_{n\times n}$. For computational efficiency, we use a small real value for $\lambda$ rather than taking the limit, which places an upper bound on the minimum error covariance equal to $1/\lambda$. Thus, $\lambda$ must be chosen carefully so as not to inject too much “artificial” observability into the system while still avoiding numerical instabilities associated with computing a numerical inverse. We set $\lambda=$1\text{\times}{10}^{-6}\text{\,}$\$ for all calculations.
6. Perform coordinate transformation
The coordinates of both $\mathcal{O}$ and $\mathcal{F}$ are principally defined in the state space $\mathcal{X}$, which means that $\mathcal{F}^{-1}$ only provides information with respect to the individual state variables in $\mathbf{x}$. However, in many cases, it is beneficial to evaluate observability in a set of new coordinates $\mathbf{z}$. For instance, a system may be defined in Cartesian coordinates, but the polar representation of the state may be the most relevant for observability analysis. The system itself could be redefined in the new coordinates $\mathbf{f}(\mathbf{z}_{k},\mathbf{u}_{k})$ and the observability analysis pipeline could be repeated. However, this process is often tedious for complex systems, not computationally efficient, and is not practical for numerical solvers that rely on a specific coordinate frame. We incorporate a simple method to transform the coordinates of $\mathcal{O}$ (or $\mathcal{F}$) from an arbitrary state space coordinate basis $\mathbf{x}$ to a new basis $\mathbf{z}$ that bypasses these drawbacks and can be appended post hoc after computing $\mathcal{O}$ with respect to $\mathbf{x}$.
First, the new basis $\mathbf{z}$ is defined as a transformation of $\mathbf{x}$ such that $\mathbf{z}=T(\mathbf{x})$, where $T(\mathbf{x})$ is a diffeomorphism that preserves the original information content in $\mathbf{x}$ 61. Because $\mathcal{O}$ fundamentally represents the Jacobian ${\partial\mathbf{Y}}/{\partial\mathbf{x}}$ in continuous time or ${\Delta\mathbf{Y}}/{\Delta\mathbf{x}}$ in discrete time, we can compute a coordinate transformation as the Jacobian of $\mathbf{x}$ with respect to $\mathbf{z}$, evaluated at the initial state $\mathbf{z}_{0}=T\mathbf{x}_{0}$
$\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{z}=\mathbf{z}_{0}}=\left.\left(\frac{\partial T(\mathbf{x})}{\partial\mathbf{x}}\right)^{-1}\right|_{\mathbf{x}=\mathbf{x}_{0}}$ (14)
and transform the coordinates as
$\mathcal{O}_{\mathbf{z}}=\left.\frac{\partial\mathbf{Y}}{\partial\mathbf{z}}\right|_{\mathbf{z}=\mathbf{z}_{0}}=\frac{\partial\mathbf{Y}}{\partial\mathbf{x}}\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{z}=\mathbf{z}_{0}}=\mathcal{O}_{\mathbf{x}}\left.\frac{\partial\mathbf{x}}{\partial\mathbf{z}}\right|_{\mathbf{z}=\mathbf{z}_{0}}.$ (15)
In practice, $\partial\mathbf{x}/\partial\mathbf{z}$ can be computed numerically or symbolically and evaluated at the initial state $\mathbf{z}_{0}$. In some cases, it may be necessary to perform multiple coordinate transforms to evaluate the observability of new state variables that may not jointly satisfy the diffeomorphism/information requirement in a single state space.
Additional considerations for observability calculations
Interpreting minimum error variance for circular variables
The minimum error variance (MEV) provided by BOUNDS requires careful interpretation for circular variables. The value returned by BOUNDS is a variance measure that assumes that the underlying variable is linear. To correctly interpret this value for a circular variable, we can transform it into a circular variance that corresponds to a wrapped normal distribution with a standard deviation equal to $\sqrt{MEV}$. This circular variance can be found according to the following equation:
$\operatorname{Var_{circ}}=1-\exp{(-MEV/2)},$ (16)
where $MEV$ is the minimum error variance given BOUNDS, e.g. $\mathcal{F}_{i}^{-1}$.
Observability for non-contiguous time windows
In some cases it is important to consider the times at which measurements are collected when evaluating active sensing motifs. For some sets of measurements, information must be collected continuously throughout a motif to increase observability (LABEL:fig:figure_supp_timestep_a). However, other sets of measurements may allow information to be considered only before and after the motif, with a similar increase in observability to continuously collecting information (LABEL:fig:figure_supp_timestep_b). This is an especially important consideration when dealing with systems where sensor noise may be correlated with a motif. For instance, translational optic flow cues may be noisier or less reliable when turning due to the influence of rotational optic flow.